# The Quantum Measurement Problem: A Review of Recent Trends

### v1

@@ 5.4.1 Physical Restrictions on Hidden Variables
The hidden variable theories are usually framed within the limits of two no-go theorems and a couple of physically motivated restrictions. The theorems are the Bell and the Kochen-Specker ones. The Bell’s theorem states that no local hidden-variable theory can be entirely compatible with quantum mechanics [87, Ch 2] [88]. Local in this context means that “the result of a measurement on one system be unaffected by operations on a distant system with which it has interacted in the past” [87, Ch 2].
In turn, the Kocher-Specker theorem [88, 89, 90] states that any hidden variable theory consistent with quantum mechanics must renounce either all observables having definite values at all times or these values being independent of how they are measured. The theorem is strictly valid for Hilbert spaces with dimension superior to two.
@@
The situation seems simple from there. For instance, the validity of Bell’s Theorem–which experiments have vastly supported [91]–seems to imply that physical theories must either be non-local or not have properties defined at all times. However, as mentioned before, no-go theorems are as strong as their underlying hypotheses, often deeply hidden in the deduction. The restrictions imposed by these theorems and physical principles have been constantly challenged [92, 93] (except for the non-signaling principle, which is likely the only consensual rule in the game). Relaxing statistical independence (sometimes loosely called the free-will principle) in Bell’s theorem, for example, may allow for building consistent local and realist theories. We will return to this point in Section 5.5.
@@ 5.4.2 Bohmian Mechanics
Most of the practical development in hidden-variable theories happens in the frame of Bohmian Mechanics, which poses the existence of corpuscules moving in a non-Newtonian way and guided by a wave function [94]. The corpuscule’s velocities are determined by a velocity field (the de Broglie-Bohm Equation) dependent on the wave function, which, in turn, follows the Schrödinger evolution. This deterministic time evolution in Bohmian Mechanics automatically solves the problem of outcomes. Although Bohmian Mechanics’ experimental predictions are equivalent to quantum mechanics [95], it is not discarded that they could be experimentally tested [96].
Poirier has proposed an alternative theory prescribing non-Newtonian trajectories but entirely discarding wave functions [97, 98]. In this case, the quantum system is described by an ensemble of interacting deterministic real-valued trajectories with well-defined positions and momenta. This type of theory inscribes itself into the tradition of hydrodynamic interpretation of quantum mechanics, inaugurated by Madelung right after Schrödinger’s original work [99][100, p. 222]. In Madelung’s interpretation, quantum mechanics is modeled as a quantum fluid [101], which can be understood as many interacting classical worlds [102].
This picture of many interacting classical worlds has been the basis of molecular dynamics simulations, where the nuclear wave function time evolution is commonly approximated by interacting and non-interacting trajectories, depending on the approach [103, 104]. Tully, who proposed surface hopping, the most popular method for nonadiabatic dynamics simulations, employed the Madelung Hydrodynamic Interpretation to justify using mixed quantum-classical approximations [105]. These molecular dynamics simulations, however, adopt the trajectory approach pragmatically without any foundational claim.
@@ 5.5 Deterministic Models
Bell’s theorem shows no local hidden variable theories can reproduce all quantum mechanical predictions [87]. Thus, the experimental violation of Bell’s inequalities seemed to be the end of any local deterministic model. However, Bell’s theorem explicitly requires the validity of statistical independence assumption, which states that the experimental settings are not influenced by any hidden factors determining the system’s evolution. This section surveys two local deterministic models—Superdeterminism and Retrocausality—built on the challenge of the statistical independence assumption.
@@
Recent Cosmic Bell Experiments have provided stringent tests of the statistical independence assumption in Bell’s theorem. These experiments avoid the possibility that hidden variables could influence detector settings by randomly determining measurement choices using distant astrophysical sources. By doing so, they attempt to close the freedom-of-choice loophole, which is crucial for evaluating the plausibility of superdeterministic explanations. In a landmark experiment, Rauch et al. [91] used high-redshift quasars—whose light originated billions of years ago—to set measurement parameters in a Bell test. Their results remained consistent with standard quantum mechanics, implying that superdeterministic correlations, if they exist, should date back to the early universe. Similarly, Handsteiner et al. [108] employed light from distant galaxies to eliminate potential local influences on detector choices. While these results do not strictly rule out superdeterminism, to remain viable, it must account for cosmological-scale correlations, raising deep challenges for its plausibility.
@@
Retrocausality is still an active research field. Adlam [112] proposed that accepting nonlocality and imposing relativistic constraints (such as the nonexistence of a preferred reference frame) leads to retrocausality. Sutherland [113] derives the Born rule within a retrocausal model. This is a significant result because it changes the status of the rule from a postulate to a natural consequence of incorporating time-symmetric boundary conditions in quantum mechanics. A notable recent development is Ridley and Adlam’s Fixed-Point Formulation [114], which presents an atemporal, all-at-once framework for retrocausality.
@@ 5.6.1 Classical-Apparatus Inducing Collapse
The core of the Copenhagen Interpretation of quantum mechanics, developed by Bohr and Heisenberg, is the assumption that any experiment in physics must be described in terms of classical physical concepts [32, Introduction]. When such an experiment is in the realm of quantum mechanics, the tension between the system’s quantum description and the apparatus’s classical description is the crucial factor causing statistical uncertainty and ultimately selecting a single output [116]. There is no claim that the classical apparatus is not composed of quantum particles. It is just that the description of the experiment is restricted to the use of classical terminology.
This quantum-classical duality, formulated as the Heisenberg Cut arbitrarily separating classical from quantum systems, still plays a role in a pragmatic definition of measurement [117]. It is also at the basis of the asymptotic emergence hypothesis, which poses that collapse is a unitary physical process. Asymptotic emergence employs algebraic quantum theory to demonstrate that some phenomena forbidden in the quantum limit (like collapse) are allowed in the classical limit due to some exponential sensitivity to perturbations in the Hamiltonian [118].
In a series of works [119, 120], Schonfeld has explored cloud chambers, Geiger counters, and Stern-Gerlach experiments for empirical tests of quantum measurement, including possible deviations from the Born rule in detecting rare events. He challenges the conventional view of measurement as an axiomatic feature of quantum mechanics, proposing that measurement emerges phenomenologically from collective interactions in quantum detection systems.
@@ 5.6.2 Mind-Inducing Collapse
Then, there is an epistemic proposal where the mind doesn’t objectively cause the collapse. Still, the collapse is just an illusion caused by the observer becoming aware of the measurement outcome. We find such perspective in London and Bauer [123, p. 251] who, in 1939, attributed the essential role played by consciousness to be “the increase of knowledge, acquired by the observation.” This is essentially the same hypothesis that sustains QBism nowadays, as discussed in Section 5.2.
@@ 5.7 Measurement in Quantum Field Theory
Quantum Field Theory (QFT) poses specific challenges to the Measurement Problem that are more daunting than in standard, non-relativistic quantum mechanics [125]. Because QFT integrates quantum mechanics and special relativity, locality is a built-in concept. This differs entirely from non-relativistic quantum mechanics, where locality is an add-on assumption, like when Bell imposed it in deriving his theorem. Sorkin showed that if one attempts to apply standard collapse theory in QFT, it may lead to superluminal information exchange in what he denoted impossible measurements [126, 127].
@@
Figure 8 illustrates a version of Sorkin’s impossible measurement using our NH3 qubit. Alice, Bob, and Charlie are ready to perform measurements on a molecule-field coupled system at times (Alice), (Bob), and (Charlie). They are spaced so that causal effects are allowed between Bob and Alice and between Charlie and Bob but not between Alice and Charlie. In the relativity lingo, Alice and Charlie are spacelike separated.
@@
Alice prepares a molecule in a state and can choose if she will flip it to at time or not. The molecule is coupled to an electromagnetic field. The field state is if the molecule is and if the molecule is . The field is initially and, after , it is updated according to Alice’s choice. At , Bob measures the field projected on . Charlie, who, at time , has no direct information about Alice’s choice, attempts to indirectly discover that by measuring the field at , which was updated after Bob’s intervention. Sorkins argues that, in principle, Charlie’s measurements on the field could reveal Alice’s choice, implying superluminal exchange. Therefore, avoiding such impossible measurements would require nontrivial constraints on which observables are consistent with causality.
@@
Such imprecision in the description of the measurement in QFT has been characterized, with a flair to the dramatic, as “a major scandal in the foundations of quantum physics” [128]. Recently, Bostelmann, Fewster, and Ruep proposed that the situation is naturally solved within algebraic QFT as long as probes and couplings are constrained to be local [125, 129]. Their analysis is valid for general fields and spacetime geometries, making it a promising step toward a rigorous measurement theory in QFT. On the other hand, Polo-Gómez, Garay, and Martín-Martínez propose that a measurement theory in QFT could be built by exploring spatially smeared detector models [130].
@@ 6 The Ontological Status of the Wave function
Other -ontic interpretations have been developed to avoid the conceptual challenges of high-dimensional spaces. One approach is the multi-field interpretation, which states that an -body wave function is a muti-field that assigns properties to sets of points in the 3D space [135]. This is the view one of us adopted in an essay discussing the nature of the molecular wave function [136]. Gao [137] proposed an alternative -ontic interpretation in the 3D space that does not require multi-fields. In his view, the wave function represents a random discontinuous motion (RDM) of particles in real space. The wave function phase encodes momentum flow, giving it a concrete physical meaning. Moreover, RDM makes testable predictions, suggesting that collapse is an emergent effect of underlying stochastic motion rather than a fundamental postulate. Whether this model extends to relativistic Quantum Field Theory remains an open question.
@@
A different approach comes from Nomological Interpretations, which suggest that the wave function is not a physical object but a law of nature governing quantum behavior [132]. This is a common stance in the interpretation of Bohmian Mechanics [94, 140].
The Berry phase, which may arise during the cyclic evolution of quantum systems [141, 142], is a mathematical feature of the wave function that could be explored to assess its ontological status. Such a phase can build, for instance, when the nuclear geometry of a molecule evolves along a closed path encircling a conical intersection (a point of electronic state degeneracy). Unlike dynamic phases, the Berry phase is tied to the wave function’s geometric properties and may, in principle, manifest in observable interference effects. Valahu et al. [143] claimed to have observed such geometric-phase interference in a trapped-ion quantum simulator mimicking conical intersection dynamics, which would strongly support the -ontic interpretation. However, it’s unclear if such effects could be observed in actual molecules since their experiment could be considered more a controlled simulation than a molecular measurement. Moreover, Min et al. [144] and, more recently, Ibele et al. [145] have provided strong theoretical evidence that the Berry phase originates from the Born-Oppenheimer approximation and disappears in an exact electron-nuclear treatment, suggesting it may be an artifact rather than a fundamental quantum property.
Although the ontological nature of the wave function remains unsettled, recent experiments provided empirical support for wave function realism. Costello et al. [146] reconstructed Bloch wave functions in GaAs, demonstrating that wave functions are not just inferred but can be directly measured. Their work highlights that, at least in solid-state physics, the wave function has an observable physical structure. This challenges purely -epistemic interpretations and suggests that wave functions, at least in some contexts, correspond to real physical entities. Nevertheless, the scope of their findings is limited because the wave functions in the study describe quasiparticles (holes in semiconductors), they are not directly the same as the universal wave function of quantum mechanics.
@@ 7 Conclusion
It is easy to get lost in the Quantum Measurement Problem. The gargantuan amount of literature, the innumerable distinct theories, the interminable debates, and the entanglement between physics, philosophy, and formal logic all lend anyone stepping on this field a feeling analogous to being trapped in a dense tropical forest. This review intends to be an updated global map, which, although it cannot point out the escape route, as it simply doesn’t exist, will at least help us stay far from dangerous cliffs.
@@
Most scientists who use quantum mechanics in their daily work focus on its practical applications rather than its interpretation. Indeed, it makes no difference if we interpret the result of an experiment as the quantum state collapsing into a single branch or experiencing a branch that cuts communication from all others. Such subtle distinction belongs to defining a preferred worldview. It is even less relevant than choosing between Newtonian or Hamiltonian descriptions of classical dynamics, which have at least instrumental implications. The exception concerns objective collapse theories, which propose new physics beyond conventional quantum mechanics. In any case, it makes sense to keep challenging all interpretations and theories to prune the weak and reach complete consistency on the remaining, even if we never get a consensual single description of quantum mechanics.

### v2

@@ 5.4.1 Physical Restrictions on Hidden Variables
The hidden variable theories are usually framed within the limits of two no-go theorems and a couple of physically motivated restrictions. The theorems are the Bell and the Kochen-Specker ones. Bell’s theorem states that no local hidden-variable theory can be entirely compatible with quantum mechanics [104, Ch 2] [105]. Local in this context means that “the result of a measurement on one system be unaffected by operations on a distant system with which it has interacted in the past” [104, Ch 2].
In turn, the Kochen-Specker theorem [105, 106, 107] states that any hidden variable theory consistent with quantum mechanics must renounce either all observables having definite values at all times or these values being independent of how they are measured. The theorem is strictly valid for Hilbert spaces with dimension greater than two.
@@
The situation seems simple from there. For instance, the validity of Bell’s Theorem—which experiments have vastly supported [108]—seems to imply that physical theories must either be non-local or not have properties defined at all times. However, no-go theorems are as strong as their underlying hypotheses, often deeply hidden in the deduction. The restrictions imposed by these theorems and physical principles have been constantly challenged [109, 110] (except for the non-signaling principle, which is likely the only consensual rule in the game). Relaxing statistical independence (sometimes loosely called the free-will principle) in Bell’s theorem, for example, may allow for building consistent local and realist theories. We will return to this point in Section 5.5.
@@ 5.4.2 Bohmian Mechanics
Most of the practical development in hidden-variable theories happens in the framework of Bohmian Mechanics, which posits the existence of corpuscules moving in a non-Newtonian way and guided by a wave function [111]. The corpuscule’s velocities are determined by a velocity field (the de Broglie-Bohm Equation) dependent on the wave function, which, in turn, follows the Schrödinger evolution. This deterministic time evolution in Bohmian Mechanics automatically solves the problem of outcomes. Although Bohmian Mechanics’ experimental predictions are equivalent to quantum mechanics [112], it is not discarded that they could be experimentally tested [113].
Poirier has proposed an alternative theory that prescribes non-Newtonian trajectories but entirely discards wave functions [114, 115]. In this case, the quantum system is described by an ensemble of interacting deterministic real-valued trajectories with well-defined positions and momenta. This type of theory inscribes itself into the tradition of hydrodynamic interpretation of quantum mechanics, inaugurated by Madelung right after Schrödinger’s original work [116][117, p. 222]. In Madelung’s interpretation, quantum mechanics is modeled as a quantum fluid [118], which can be understood as many interacting classical worlds [119].
This picture of many interacting classical worlds has been the basis of molecular dynamics simulations, where the nuclear wave function time evolution is commonly approximated by interacting and non-interacting trajectories, depending on the approach [120, 121]. Tully, who proposed surface hopping, the most popular method for nonadiabatic molecular dynamics simulations, employed the Madelung Hydrodynamic Interpretation to justify using mixed quantum-classical approximations [122]. These molecular dynamics simulations, however, adopt the trajectory approach pragmatically without any foundational claim.
@@ 5.5 Deterministic Models
Bell’s theorem shows that no local hidden variable theories can reproduce all quantum mechanical predictions [104]. Thus, the experimental violation of Bell’s inequalities seemed to be the end of any local deterministic model. However, Bell’s theorem explicitly requires the validity of statistical independence assumption, which states that the experimental settings are not influenced by any hidden factors determining the system’s evolution. This section surveys two local deterministic models—Superdeterminism and Retrocausality—built on the challenge of the statistical independence assumption.
@@
Cosmic Bell Experiments have provided stringent tests of the statistical independence assumption in Bell’s theorem [108, 125]. These experiments avoid the possibility that hidden variables could influence detector settings by randomly determining measurement choices using distant astrophysical sources. By doing so, they attempt to close the freedom-of-choice loophole, which is crucial for evaluating the plausibility of superdeterministic explanations. In a landmark experiment, Rauch et al. [108] used high-redshift quasars—whose light originated billions of years ago—to set measurement parameters in a Bell test. Their results remained consistent with standard quantum mechanics, implying that superdeterministic correlations, if they exist, should date back to the early universe. Similarly, Handsteiner et al. [125] employed light from distant galaxies to eliminate potential local influences on detector choices. While these results do not strictly rule out superdeterminism, it must account for cosmological-scale correlations to remain viable, raising deep challenges for its plausibility.
@@
Retrocausality is still an active research field. Adlam [129] proposed that accepting nonlocality and imposing relativistic constraints (such as the nonexistence of a preferred reference frame) leads to retrocausality. Sutherland [130] derives the Born rule within a retrocausal model. This is a significant result because it changes the status of the rule from a postulate to a natural consequence of incorporating time-symmetric boundary conditions in quantum mechanics. A notable development is the Fixed-Point Formulation proposed by Ridley and Adlam, which presents an atemporal, all-at-once framework for retrocausality and includes a full derivation of the Born rule [131, 132]. Another recent contribution is van der Pals’ retrocausal hidden-variable model [133], which attributes the emergence of definite outcomes to resonance conditions between advanced and retarded virtual photons. These photons arise from an underlying, time-symmetric periodic process occurring beneath the Heisenberg uncertainty threshold.
@@ 5.6.1 Classical-Apparatus Inducing Collapse
The core of the Copenhagen Interpretation of quantum mechanics, developed by Bohr and Heisenberg, is the assumption that any experiment in physics must be described in terms of classical physical concepts [39, Introduction]. When such an experiment is in the realm of quantum mechanics, the tension between the system’s quantum description and the apparatus’s classical description is the crucial factor causing statistical uncertainty and ultimately selecting a single output [135]. There is no claim that the classical apparatus is not composed of quantum particles. It is just that the description of the experiment is restricted to the use of classical terminology.
This quantum-classical duality, formulated as the Heisenberg Cut arbitrarily separating classical from quantum systems, still plays a role in a pragmatic definition of measurement [136]. It is also at the basis of the asymptotic emergence hypothesis [137], which poses that collapse is a unitary physical process. Asymptotic emergence employs algebraic quantum theory to demonstrate that some phenomena forbidden in the quantum limit (like collapse) are allowed in the classical limit due to some exponential sensitivity to perturbations in the Hamiltonian [137].
In a series of works [138, 139], Schonfeld has explored cloud chambers, Geiger counters, and Stern-Gerlach experiments for empirical tests of quantum measurement, including possible deviations from the Born rule in detecting rare events. He challenges the conventional view of measurement as an axiomatic feature of quantum mechanics, proposing that measurement emerges phenomenologically from collective interactions in quantum detection systems. He analyzes the detailed microstructure of real measurement systems, using classical idealizations selectively where it seems mathematically or intuitively reasonable.
@@ 5.6.2 Mind-Inducing Collapse
Then, there is an epistemic proposal where the mind doesn’t objectively cause the collapse. Still, the collapse is just an illusion caused by the observer becoming aware of the measurement outcome. We find such a perspective in London and Bauer [73, p. 251] who, in 1939, attributed the essential role played by consciousness to be “the increase of knowledge, acquired by the observation.”
@@ 5.7 Measurement in Quantum Field Theory
Quantum Field Theory (QFT) poses specific challenges to the Measurement Problem that are more daunting than in standard, non-relativistic quantum mechanics [143]. Because QFT integrates quantum mechanics and special relativity, locality is a built-in concept. This differs entirely from non-relativistic quantum mechanics, where locality is an add-on assumption, like when Bell imposed it in deriving his theorem. Sorkin showed that applying standard collapse rules to QFT can lead to superluminal information transfer in scenarios he termed impossible measurements [144]. More recently, this issue has been reframed as a violation of the broader no-signaling principle, which prohibits communication without a physical carrier [145].
@@
Figure 8 illustrates a version of Sorkin’s impossible measurement using our NH3 qubit. Alice, Bob, and Charlie are ready to perform measurements on a molecule-field coupled system at times (Alice), (Bob), and (Charlie). They are spaced to allow causal effects between Bob and Alice and between Charlie and Bob, but not between Alice and Charlie. In the relativity lingo, Alice and Charlie are spacelike separated.
@@
Alice prepares a molecule in a state and can choose whether to flip it to at time or not. The molecule is coupled to an electromagnetic field. The field state is if the molecule is and if the molecule is . The field is initially and, after , it is updated according to Alice’s choice. At , Bob measures the field projected on . Charlie, who, at time , has no direct information about Alice’s choice, attempts to indirectly discover that by measuring the field at , which was updated after Bob’s intervention. Sorkin argues that, in principle, Charlie’s measurements on the field could reveal Alice’s choice, implying superluminal exchange. Therefore, avoiding such impossible measurements would require nontrivial constraints on which observables are consistent with causality.
@@
Such imprecision in the description of the measurement in QFT has been characterized, with a flair to the dramatic, as “a major scandal in the foundations of quantum physics” [146]. Recently, Bostelmann, Fewster, and Ruep proposed that the situation is naturally solved within algebraic QFT as long as probes and couplings are constrained to be local [147, 143, 148].
This approach, originally developed as a general framework for describing measurement in algebraic QFT, was not tailored to solve the Sorkin problem, but rather aims to identify and localize induced system observables and to characterize state updates. The application to the Sorkin scenario, as presented in [148], thus constitutes a nontrivial test of the framework. Their analysis is valid for general fields and spacetime geometries, making it a promising step toward a rigorous measurement theory in QFT.
A complementary approach is proposed by Polo-Gómez, Garay, and Martín-Martínez, who develop a measurement theory for QFT based on spatially smeared particle detector models—such as the Unruh-DeWitt detector—that interact locally and covariantly with quantum fields [149]. Their framework, applicable to general fields and spacetime geometries, avoids problems like Sorkin’s impossible measurements and allows localized interactions without invoking instantaneous collapse. While Bostelmann et al. pursue a mathematically rigorous algebraic route, Polo-Gómez et al. adopt a more operational strategy grounded in how detectors realistically interact with quantum fields.
Other proposals inspired by quantum field theory and signal analysis take a different route, using mathematical tools to reinterpret collapse as a statistical construction. For instance, Morgan proposed a framework based on classical random fields and operator-algebraic methods to model quantum phenomena, offering an alternative strategy that remains within a classical probabilistic setting [150]. In his work, these fields—equipped with nontrivial correlation structures—are used to reproduce quantum statistics and construct joint probabilities even for noncommuting observables. His emphasis lies not on deriving definite outcomes from quantum theory, but on modeling the empirical structure of datasets generated by repeated measurements. This approach aims to preserve locality and realism, challenging standard assumptions about the necessity of wave function collapse or quantization itself.
@@ 6 The Ontological Status of the Wave Function
Other -ontic interpretations have been developed to avoid the conceptual challenges of high-dimensional spaces. One approach is the multi-field interpretation, which states that an -body wave function is a multi-field that assigns properties to sets of points in the 3D space [155]. This is the view one of us adopted in an essay discussing the nature of the molecular wave function [156]. Gao [157] proposed an alternative -ontic interpretation in the 3D space that does not require multi-fields. In his view, the wave function represents a random discontinuous motion (RDM) of particles in real space. The wave function phase encodes momentum flow, giving it a concrete physical meaning. Moreover, RDM makes testable predictions, suggesting that collapse is an emergent effect of underlying stochastic motion rather than a fundamental postulate. Whether this model extends to relativistic Quantum Field Theory remains an open question.
@@
A different approach comes from Nomological Interpretations, which suggest that the wave function is not a physical object but a law of nature governing quantum behavior [152]. This is a common stance in interpreting Bohmian Mechanics [111, 160].
The Berry phase, which may arise during the cyclic evolution of quantum systems [161, 162], is a mathematical feature of the wave function that could be explored to assess its ontological status. Such a phase can build, for instance, when the nuclear geometry of a molecule evolves along a closed path encircling a conical intersection. (A conical intersection is a molecular geometry at which two adiabatic electronic states degenerate, forming a conical topology of the potential energy surfaces in nuclear coordinate space [163].) Unlike dynamic phases, the Berry phase is tied to the wave function’s geometric properties and may, in principle, manifest in observable interference effects. Valahu et al. [164] claimed to have observed such geometric-phase interference in a trapped-ion quantum simulator mimicking conical intersection dynamics, which would strongly support the -ontic interpretation. However, it’s unclear if such effects could be observed in actual molecules since their experiment could be considered more a controlled simulation than a molecular measurement. Moreover, Min et al. [165] and, more recently, Ibele et al. [166] have provided strong theoretical evidence that the Berry phase originates from the Born-Oppenheimer approximation and disappears in an exact electron-nuclear treatment, suggesting it may be an artifact rather than a fundamental quantum property.
Although the wave function’s ontological nature remains unsettled, some recent experiments may support wave function realism. In particular, Costello et al. [167] reconstructed Bloch wave functions in GaAs semiconductors using angle-resolved photoemission data, showing that wave functions can, in some contexts, be experimentally accessed rather than only inferred. However, such reconstructions apply to quasiparticles and rely on a specific preparation context, and do not settle the deeper question of whether the wave function represents an element of reality. As Rovelli has emphasized [69], a wave function may encode the full information needed to predict measurement outcomes relative to a given observer, without implying a universally ontic status.
@@ 7 Conclusion
It is easy to get lost in the Quantum Measurement Problem. The gargantuan amount of literature, the innumerable distinct theories, the interminable debates, and the entanglement between physics, philosophy, and formal logic all lend anyone stepping into this field a feeling analogous to being trapped in a dense tropical forest. This review intends to be an updated global map, which, although it cannot point out the escape route, as it simply doesn’t exist, will at least help us stay far from dangerous cliffs.
@@
Most scientists who use quantum mechanics in their daily work focus on its practical applications rather than its interpretation. Indeed, it makes no difference if we interpret the result of an experiment as the quantum state collapsing into a single branch or experiencing a branch that cuts communication from all others. Such a subtle distinction belongs to defining a preferred worldview. It is even less relevant than choosing between Newtonian or Hamiltonian descriptions of classical dynamics, which have at least instrumental implications. The exception concerns objective collapse theories, which propose new physics beyond conventional quantum mechanics. In any case, it makes sense to keep challenging all interpretations and theories to prune the weak and reach complete consistency on the remaining, even if we never get a consensual single description of quantum mechanics.